# Quantum tunneling and defect-induced transport modulation in twisted bilayer graphene superlattices

### v1

@@ I INTRODUCTION
Beyond monolayer graphene, graphene superlattices provide a useful route to control the electronic band structure and the transport properties of Dirac quasiparticles. In particular, theoretical and experimental studies on electrostatic and magnetic graphene superlattices have demonstrated tunable transmission gaps and rich tunneling behaviors, enabling practical device applications like wave-vector filters and tunneling magnetoresistance components [10.1063/1.4729133, 10.1063/1.3626278, PhysRevB.79.165412, 10.1063/1.3467778, PhysRevB.77.115446, 10.1063/1.4953865]. In particular, finite graphene superlattices exhibit a universal resonance-splitting rule, where each miniband splits into (N1)(N–1) resonance peaks for a structure with $N$ barriers, analogous to semiconductor heterostructures first described by Tsu and Esaki [10.1063/1.1654509, Pham_2015, luo2022filtering]. Many studies have extended this concept, showing that miniband formation and resonant tunneling can be tuned by external fields and structural modulation [Tran_2024]. These advances motivate our focus on twist-tunable superlattices in TBG, where structural control (twist angle) can be combined with superlattice design to engineer transmission windows and angular selectivity for device applications.
@@ II MODEL
where $E+\Delta U_{0}$ and $E+\Delta U_{d}$ are the barrier height and defect height, respectively. $\Delta U_{0}$ ($\Delta U_{d}$) is the energy difference between the height barriers (defect height) and the incident energy $E$. $d_{B}$ and $d_{w}$ are, respectively, the barrier and well widths. The distance $d=d_{B}+d_{w}$ represents the width of the unit cell, which is composed of one barrier and one well. The $j$ is an integer, $1\leq j\leq N$, with $N$ potential barriers. The $j_{d}$ is the index of the defect barrier. We assume the $N$ potential barriers with a rectangular shape, which prevents neglecting intervalley scattering between $K$ and $K_{\theta}$ [he2013chiral, katsnelson2006chiral]. Therefore, the scattering process is analyzed within a single valley.
@@ III Results and discussion
First, we analyzed the behavior of the transmission probability of normally incident electrons as a function of incident energy $E$, as shown in Fig. 2(a)-(f). Here, the superlattice structure is symmetric and the number of barriers is varied between $N=4$ and $9$. In this case, we only consider the effect of an electric potential with a twist angle of $\theta=3.89^{\circ}$. Parameters used are: the height of barriers $\Delta U_{0}=0.3eV$, the width of barriers $d_{B}=30\;nm$, and wells $d_{w}=25\;nm$. As we can see, the transmission profile shows oscillations for all cases. These oscillations increase as the number of barriers increases. The transmission coefficient exhibits four series of resonances separated by three minigaps, and these series depend on the number of barriers. Each series of resonances contains N1N-1 resonant states. This feature, first demonstrated for semiconductor superlattices by Tsu et al. [10.1063/1.1654509] and later established for graphene-based finite superlattices [xu2014resonant, Pham_2015], arises from the splitting of quasi-bound states into coupled minibands as the number of barriers increases. The resonant peaks move closer to each other and become sharper as the number of barriers increases. Furthermore, we observe that the depths of minigaps depend on the number of barriers; they become more pronounced and deeper for higher values of $N$. For a large number of barriers, we can see that the transmission, particularly in the third minigap, is completely suppressed, resulting in a transmission gap due to the existence of evanescent modes inside the superlattice.
@@
To further understand the dependence of the incident angle on the tunneling of incident electrons through the structure, we examine the variation of the transmission coefficient as a function of incident angle $T\left(\varphi\right)$ for different $N$ and twist angles $\theta$, as shown in Fig. 5. For low incident energy and a small value of $\Delta U_{0}$, the angular dependence of the transmission probability for the TBG resembles that of a graphene monolayer superlattices [PhysRevB.76.075430], where the normally incident electrons tunnel perfectly or almost perfectly, regardless of the twist angle $\theta$ and the number of barriers $N$ (see the upper panels of Fig. 5). However, important differences emerge for large incident energy and large value of $\Delta U_{0}$. From the lower panel of Fig. 5, we see that for the TBG configuration with a twist angle of $\theta=3.89^{\circ}$, $T\left(\varphi\right)$ shows that the chiral fermions tunnel perfectly or almost perfectly at normal incidence. However, the asymmetry of the transmission about $\varphi=0^{\circ}$ increases with increasing barrier number $N$. In addition, more resonant tunneling peaks appear as a function of the incident angle. This indicates that the number of barriers plays a significant role in the anisotropic transmission of the TBG superlattice. The perfect or almost tunneling and resonant tunneling occur for a wide range of incident angles for $\theta=3.89^{\circ}$, as can be seen from lower panel of Fig. 5. Importantly, for large $E$ and $\Delta U_{0}$, the perfect transmission of normally incident electrons depends sensitively on the twist angle $\theta$ (see Fig. 5(d)-(f)). This is due to that the separation between the two Dirac cones depends on the twist angle, and thus the chiral character or pseudospins of the quasiparticles involved.
@@
Now, we consider the case of TBG superlattices with a defect; we plot the transmission coefficient as a function of incident angle for $\Delta U_{d}>\Delta U_{0}$ and $\Delta U_{d}<\Delta U_{0}$, as shown in Fig. 6(a)-(c) and Fig. 6(d)-(f), respectively. Particularly, we observe that for the configuration with $\theta=3.89^{\circ}$, the angular behavior of $T\left(\varphi\right)$ at normal incidence is similar to that of the periodic structure without a defect, and the chiral fermions are always tunnel perfectly or almost perfectly at normal incidence. This implies that the tunneling behavior in the TBG configuration with a twist angle of $\theta=3.89^{\circ}$ shares similarities with those in monolayer graphene. However, for TBG configurations with smaller twist angles, particularly $\theta=2.16^{\circ}$, the transmission probability or normally incident electrons shows a gradual transition, and can be changed from perfect tunneling to complete reflection by controlling either the twist angle or the TBG superlattice parameters (i.e., $\Delta U_{d}$, $\Delta U_{0}$ and $N$). When applying a defect at a smaller twist angle, the transmission decreases and becomes narrower and shorter on the negative angle side, and the Klein tunneling effect turns off. For $\Delta U_{d}<\Delta U_{0}$, for a larger number of barriers, we observe that for the configuration with $\theta=2.16^{\circ}$, the incident electrons emitted from the $K$-cone are deflected towards negative angles with the suppression of Klein tunneling (Fig. 6(f)). In contrast, because the mirror-symmetric behavior of the two cones, the electrons emitted from $K_{\theta}$-cone may propagate to the other side, toward positive angles, and no Klein tunneling is expected. Normal-incidence tunneling for TBG configuration with a twist angle of $\theta=3.89^{\circ}$ becomes completely forbidden for $E\geq 2E_{V}$ [he2013chiral, BAHLAOUI2024115880]. According to our findings, controlling TBG with smaller twist angles via different barrier structures, with or without a defect, should strongly affect normal tunneling and the appearance of Klein tunneling.

### v2

@@ I INTRODUCTION
Beyond monolayer graphene, graphene superlattices provide a useful route to control the electronic band structure and the transport properties of Dirac quasiparticles. In particular, theoretical and experimental studies on electrostatic and magnetic graphene superlattices have demonstrated tunable transmission gaps and rich tunneling behaviors, enabling practical device applications like wave-vector filters and tunneling magnetoresistance components [10.1063/1.4729133, 10.1063/1.3626278, PhysRevB.79.165412, 10.1063/1.3467778, PhysRevB.77.115446, 10.1063/1.4953865]. In particular, finite graphene superlattices exhibit a universal resonance-splitting rule, where each miniband splits into (N1)(N–1) resonance peaks for a structure with NN barriers, analogous to semiconductor heterostructures first described by Tsu and Esaki [10.1063/1.1654509, Pham_2015, luo2022filtering]. Many studies have extended this concept, showing that miniband formation and resonant tunneling can be tuned by external fields and structural modulation [Tran_2024]. These advances motivate our focus on twist-tunable superlattices in TBG, where structural control (twist angle) can be combined with superlattice design to engineer transmission windows and angular selectivity for device applications.
@@ II MODEL
where $E+\Delta U_{0}$ and $E+\Delta U_{d}$ are the barrier height and defect height, respectively. $\Delta U_{0}$ ($\Delta U_{d}$) is the energy difference between the height barriers (defect height) and the incident energy $E$. $d_{B}$ and $d_{w}$ are, respectively, the barrier and well widths. The distance $d=d_{B}+d_{w}$ represents the width of the unit cell, which is composed of one barrier and one well. The $j$ is an integer, $1\leq j\leq N$, with NN potential barriers. The $j_{d}$ is the index of the defect barrier. We assume the NN potential barriers with a rectangular shape, which prevents neglecting intervalley scattering between $K$ and $K_{\theta}$ [he2013chiral, katsnelson2006chiral]. Therefore, the scattering process is analyzed within a single valley.
@@ III Results and discussion
First, we analyzed the behavior of the transmission probability of normally incident electrons as a function of incident energy $E$, as shown in Fig. 2(a)-(f). Here, the superlattice structure is symmetric and the number of barriers is varied between $N=4$ and 99. In this case, we only consider the effect of an electric potential with a twist angle of $\theta=3.89^{\circ}$. Parameters used are: the height of barriers $\Delta U_{0}=0.3eV$, the width of barriers $d_{B}=30\;nm$, and wells $d_{w}=25\;nm$. As we can see, the transmission profile shows oscillations for all cases. These oscillations increase as the number of barriers increases. The transmission coefficient exhibits four series of resonances separated by three minigaps, and these series depend on the number of barriers. Each series of resonances contains N1N-1 resonant states. This feature, first demonstrated for semiconductor superlattices by Tsu et al. [10.1063/1.1654509] and later established for graphene-based finite superlattices [xu2014resonant, Pham_2015], arises from the splitting of quasi-bound states into coupled minibands as the number of barriers increases. The resonant peaks move closer to each other and become sharper as the number of barriers increases. Furthermore, we observe that the depths of minigaps depend on the number of barriers; they become more pronounced and deeper for higher values of NN. For a large number of barriers, we can see that the transmission, particularly in the third minigap, is completely suppressed, resulting in a transmission gap due to the existence of evanescent modes inside the superlattice.
@@
To further understand the dependence of the incident angle on the tunneling of incident electrons through the structure, we examine the variation of the transmission coefficient as a function of incident angle $T\left(\varphi\right)$ for different NN and twist angles $\theta$, as shown in Fig. 5. For low incident energy and a small value of $\Delta U_{0}$, the angular dependence of the transmission probability for the TBG resembles that of a graphene monolayer superlattices [PhysRevB.76.075430], where the normally incident electrons tunnel perfectly or almost perfectly, regardless of the twist angle $\theta$ and the number of barriers NN (see the upper panels of Fig. 5). However, important differences emerge for large incident energy and large value of $\Delta U_{0}$. From the lower panel of Fig. 5, we see that for the TBG configuration with a twist angle of $\theta=3.89^{\circ}$, $T\left(\varphi\right)$ shows that the chiral fermions tunnel perfectly or almost perfectly at normal incidence. However, the asymmetry of the transmission about $\varphi=0^{\circ}$ increases with increasing barrier number NN. In addition, more resonant tunneling peaks appear as a function of the incident angle. This indicates that the number of barriers plays a significant role in the anisotropic transmission of the TBG superlattice. The perfect or almost tunneling and resonant tunneling occur for a wide range of incident angles for $\theta=3.89^{\circ}$, as can be seen from lower panel of Fig. 5. Importantly, for large $E$ and $\Delta U_{0}$, the perfect transmission of normally incident electrons depends sensitively on the twist angle $\theta$ (see Fig. 5(d)-(f)). This is due to that the separation between the two Dirac cones depends on the twist angle, and thus the chiral character or pseudospins of the quasiparticles involved.
@@
Now, we consider the case of TBG superlattices with a defect; we plot the transmission coefficient as a function of incident angle for $\Delta U_{d}>\Delta U_{0}$ and $\Delta U_{d}<\Delta U_{0}$, as shown in Fig. 6(a)-(c) and Fig. 6(d)-(f), respectively. Particularly, we observe that for the configuration with $\theta=3.89^{\circ}$, the angular behavior of $T\left(\varphi\right)$ at normal incidence is similar to that of the periodic structure without a defect, and the chiral fermions are always tunnel perfectly or almost perfectly at normal incidence. This implies that the tunneling behavior in the TBG configuration with a twist angle of $\theta=3.89^{\circ}$ shares similarities with those in monolayer graphene. However, for TBG configurations with smaller twist angles, particularly $\theta=2.16^{\circ}$, the transmission probability or normally incident electrons shows a gradual transition, and can be changed from perfect tunneling to complete reflection by controlling either the twist angle or the TBG superlattice parameters (i.e., $\Delta U_{d}$, $\Delta U_{0}$ and NN). When applying a defect at a smaller twist angle, the transmission decreases and becomes narrower and shorter on the negative angle side, and the Klein tunneling effect turns off. For $\Delta U_{d}<\Delta U_{0}$, for a larger number of barriers, we observe that for the configuration with $\theta=2.16^{\circ}$, the incident electrons emitted from the $K$-cone are deflected towards negative angles with the suppression of Klein tunneling (Fig. 6(f)). In contrast, because the mirror-symmetric behavior of the two cones, the electrons emitted from $K_{\theta}$-cone may propagate to the other side, toward positive angles, and no Klein tunneling is expected. Normal-incidence tunneling for TBG configuration with a twist angle of $\theta=3.89^{\circ}$ becomes completely forbidden for $E\geq 2E_{V}$ [he2013chiral, BAHLAOUI2024115880]. According to our findings, controlling TBG with smaller twist angles via different barrier structures, with or without a defect, should strongly affect normal tunneling and the appearance of Klein tunneling.